Solve by Elimination
x+9y+78
-x+5y=34

Answer by checkley77(12844)

x+9y+78
-x+5y=34 add
---------------
14y=112
y=112/14
y=8 ans.
x+9*8=78
x+72=78
x=78-72
x=6 ans.
Proof:
-6+5*8=34
-6+40=34
34=34